# The separating Noether number of small groups

### v2

@@ 3.8. The use of automorphisms
(i) and (ii): The subset $\rho(G)$ of $\mathrm{GL}(V)$ coincides with the subset $(\rho\circ\alpha)(G)$ of $\mathrm{GL}(V)$, therefore the representations $\rho$ and $\rho\circ\alpha$ yield the same partition of $V$ into the union of $G$-orbits, and the same polynomials in $K[V]$ are invariants (respectively relative invariants) for the $G$-action given by $\rho$ as for the $G$-action given by $\rho\circ\alpha$. Moreover, if $f\in K[(V,\rho)]^{G,\chi}$, $g\in G$ and $v\in V$, then $f((\rho\circ\alpha)(g)(v))=f(\rho(\alpha(g))(v)=\chi(\alpha(g))f(v)=(\chi\circ\alpha)(g)f(v)$. So if a relative invariant $f\in K[V]$ has weight $\chi$ with respect to $\rho$, then it has weight $\chi\circ\alpha$ with respect to $\rho\circ\alpha$. This explains both statements.
(iii): For each $g\in G$ we have
$(f\circ T)(gv)=f((T\circ\rho(g))(v))=f((\rho\circ\alpha)(g)(T(v))=\chi(\alpha(g))f(T(v))=(\chi\circ\alpha)(g)(f\circ T)(v).$
4. Groups with a cyclic subgroup of index two
Denote by $\mathrm{C}_{n}$ the cyclic group of order $n$. We use the notation for semidirect products
$\mathrm{C}_{m}\rtimes_{d}\mathrm{C}_{n}=\langle a,b\mid a^{m}=1,b^{n}=1,bab^{-1}=a^{d}\rangle\quad\text{ where }d\in\mathbb{N}\mbox{ is coprime to }m.$
We need the following classification of finite groups with a cyclic subgroup of index $2$ (see [12, Section 10] for details and references):
@@ 4. Groups with a cyclic subgroup of index two
Any finite group containing a cyclic subgroup of index two is isomorphic to
(12) $\mathrm{C}_{s}\times(\mathrm{C}_{r}\rtimes_{-1}H)$
where $r,s$ are coprime odd integers, $\mathrm{C}_{r}\rtimes_{-1}H$ is the semidirect product where the kernel of the action of $H$ on $\mathrm{C}_{r}$ contains an index two cyclic subgroup $\langle a\rangle$ of $H$, $b\in H\setminus\langle a\rangle$ acts via inversion on $\mathrm{C}_{r}$ (i.e. $bxb^{-1}=x^{-1}$ for all $x\in\mathrm{C}_{r}$), and $H$ is one of the following $2$-groups:
  • (i)
    $\mathrm{C}_{2^{n}}$  ($n\geq 1$);
  • (ii)
    $\mathrm{C}_{2^{n-1}}\times\mathrm{C}_{2}$  ($n\geq 2$);
  • (iii)
    $\mathrm{M}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{d}\mathrm{C}_{2},\qquad d={2^{n-2}+1}$  ($n\geq 3$);
  • (iv)
    $\mathrm{D}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{-1}\mathrm{C}_{2}$  ($n\geq 4$);
  • (v)
    $\mathrm{SD}_{2^{n}}:=\mathrm{C}_{2^{n-1}}\rtimes_{d}\mathrm{C}_{2},\qquad d={2^{n-2}-1}$  ($n\geq 4$);
  • (vi)
    $\mathrm{Dic}_{2^{n}}:=\langle a,b\mid a^{2^{n-1}}=1,b^{2}=a^{2^{n-2}},bab^{-1}=a^{-1}\rangle$  ($n\geq 3$).
Among the groups listed in Proposition 4.1 are the dicyclic groups. For a positive integer $m>1$, the dicyclic group is
$\mathrm{Dic}_{4m}=\begin{cases}\mathrm{C}_{r}\rtimes_{-1}\mathrm{Dic}_{2^{n}}&\text{ for }m=2^{n-2}r\text{ even }\\ \mathrm{C}_{m}\rtimes_{-1}\mathrm{C}_{4}&\text{ for }m>1\text{ odd. }\end{cases}$
Note that $\mathrm{Dic}_{8}$ is the quaternion group of order $8$.
Proposition 4.2.
Let $G$ be a non-cyclic group with a cyclic subgroup of index two. Assume that $K$ has an element of multiplicative order $|G|$. Then
$\beta_{\mathrm{sep}}^{K}(G)\geq\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
@@
Let $G$ be a group as in (12) in Proposition 4.1, where $H$ is a semidirect product of a cyclic group and the two-element group (so $H$ is $\mathrm{C}_{2}$ or $H$ is of type (ii), (iii), (iv), or (v) from Proposition 4.2). Then $G$ has a matrix representation generated by the matrices
$\ \begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\xi&0&0\\ 0&\xi^{k}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $k$ is some positive integer coprime to $2^{n-1}$ depending on the type of $H$, and $\xi$, $\rho$, $\varepsilon$ are roots of $1$ of multiplicative order $2^{n-1}$, $r$, $s$. Consider the corresponding algebra $K[x_{1},x_{2},t]^{G}$, where $x_{1},x_{2},t$ are the coordinate functions on the vector space $K^{3}$, on which the group $G$ acts via the given matrix representation. Setting $m:=sr2^{n-1}$, the monomials $x_{1}^{m}$ and $x_{2}^{m}$ are fixed by the cyclic index two subgroup $A$ of $G$ generated by the second, third, and fourth matrices above. Moreover, they are interchanged by the first matrix. It follows that $(x_{1}^{m}-x_{2}^{m})t$ is a $G$-invariant. This degree $m+1$ invariant takes different values on the vectors $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$ in $K^{3}$.
Note that our representation is the direct sum of a $2$-dimensional and a $1$-dimensional representation. So by Lemma 3.5 it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $m$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $m$. So $f=t^{d}f_{1}(x_{1},x_{2})$. Since the projections of these two vectors in the $x_{1},x_{2}$ coordinate plane are the same, no invariant depending only on $x_{1}$ and $x_{2}$ can separate them. Thus we have $d>0$. The invariants depending only on $t$ are the polynomials in $t^{2}$, which take the same value (namely $1$) on these vectors, hence $\deg(f_{1})>0$. Since $t$ is $A$-invariant, $f_{1}(x_{1},x_{2})\in K[x_{1},x_{2}]^{A}$, so $f_{1}$ is a $K$-linear combination of $A$-invariant monomials in $x_{1}$ and $x_{2}$. The monomials having positive degree in $x_{2}$ vanish on both $v$ and $v^{\prime}$. It follows that $f_{1}$ must have an $A$-invariant monomial depending only on $x_{1}$. The only such monomials are the powers of $x_{1}^{m}$. Thus $\deg(f)\geq m+d>m$, a contradiction.
Assume next that $G$ is a dicyclic group $\mathrm{Dic}_{4m}$. It is isomorphic to the matrix group generated by
$\begin{bmatrix}0&\mathrm{i}\\ \mathrm{i}&0\\ \end{bmatrix},\quad\begin{bmatrix}\rho&0\\ 0&\rho^{-1}\\ \end{bmatrix}$
where $\mathrm{i}$ is an element of $K$ with multiplicative order $4$ (so $\mathrm{i}^{2}=-1$) and $\rho$ is an element of $K$ of multiplicative order $2m$. It is an easy exercise to show that the corresponding algebra of invariants is generated by $x_{1}^{2m}-x_{2}^{2m}$, $(x_{1}x_{2})^{2}$, $(x_{1}^{2m}+x_{2}^{2m})x_{1}x_{2}$ when $m$ is odd, and by $x_{1}^{2m}+x_{2}^{2m}$, $(x_{1}x_{2})^{2}$, $(x_{1}^{2m}-x_{2}^{2m})x_{1}x_{2}$ when $m$ is even. If $m$ is odd, the vectors $[1,1]^{T}$ and $[1,-1]^{T}$ are separated by the degree $22m+$ 2 generator, but all the smaller degree generators agree on them. By assumption $K$ has an element $\nu$ of multiplicative order $4m=|G|$. For even $m$ the vectors $[\nu,1]^{T}$ and $[\nu,-1]^{T}$ are separated by the degree $22m+$ 2 generator, whereas the smaller degree generators coincide on them.
A matrix representation of $\mathrm{C}_{s}\times\mathrm{Dic}_{4m}$ (where $s>1$ is odd and is co-prime to $m$) is generated by
$\begin{bmatrix}0&\mathrm{i}&0\\ \mathrm{i}&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $\rho$ and $\varepsilon$ are roots of $1$ with multiplicative order $2m$ and $s$. The vectors $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$ are separated by the invariant $(x_{1}^{2ms}-x_{2}^{2ms})t$ when $m$ is odd and by $(x_{1}^{2ms}+x_{2}^{2ms})t$ when $m$ is even. On the other hand, any invariant of degree at most $2ms$ takes the same value on $v$ and $v^{\prime}$ (this can be seen similarly to the argument in the second paragraph of the proof, using Lemma 3.5).
It remains to deal with the case when $G\cong\mathrm{C}_{s}\times(\mathrm{C}_{r}\rtimes_{-1}\mathrm{C}_{2^{n}})$ where $n\geq 3$. A matrix representation of this group is generated by
$\begin{bmatrix}0&\omega&0\\ \omega&0&0\\ 0&0&-1\end{bmatrix},\quad\begin{bmatrix}\rho&0&0\\ 0&\rho^{-1}&0\\ 0&0&1\end{bmatrix},\quad\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&1\end{bmatrix},$
where $\omega$, $\rho$ and $\varepsilon$ are roots of $1$ with multiplicative order $2^{n}$, $r$, and $s$. The square of the first matrix together with the second and the third generate a cyclic subgroup $A$ of $G$ of order $m:=2^{n-1}rs$, acting on the line spanned by $[1,0,0]^{T}$ via a character of order $m$. The invariant $(x_{1}^{m}+x_{2}^{m})t$ separates $v:=[1,0,1]^{T}$ and $v^{\prime}:=[1,0,-1]^{T}$, and any invariant of degree at most $m$ takes the same value on $v$ and $v^{\prime}$ (this can be seen similarly to the argument in the second paragraph of the proof, using Lemma 3.5). ∎
Proof of Theorem 2.1. The Noether number $\beta^{K}(G)$ was computed for all groups with a cyclic subgroup of index $2$ in [12, Theorem 10.3], where (as a special case of a statement on the so-called generalized Noether number) it was shown that if $G$ is a non-cyclic group with a cyclic subgroup of index two and $\mathrm{char}(K)$ is not a divisor of $|G|$, then
(13) $\beta^{K}(G)=\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
Therefore the desired result on the separating Nother number is an immediate consequence of Proposition 4.2, the obvious inequality $\beta_{\mathrm{sep}}^{K}(G)\leq\beta^{K}(G)$, and (13). $\square$
4.1. The groups $\mathrm{D}_{2n}\times\mathrm{C}_{2}$
In this section $n\geq 2$ is even, and
$G=\mathrm{D}_{2n}\times\mathrm{C}_{2}=\langle a,b,c\mid a^{n}=b^{2}=c^{2}=1,\ bab=a^{-1},\ ac=ca,\ bc=cb\rangle.$
Proposition 4.3.
Assume that $K$ contains an element of multiplicative order $n$. Then we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq n+2$.
Proof.
Let $\rho$ denote an element of $K$ of multiplicative order $n$. Take the direct sum of the representations
$a\mapsto\begin{bmatrix}\rho&0\\ 0&\rho^{-1}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},$
$a\mapsto 1,\quad b\mapsto-1,\quad c\mapsto-1\qquad\text{ and }\qquad a\mapsto 1,\quad b\mapsto 1,\quad c\mapsto-1.$
Recall that $x_{1},x_{2}$ stand for the standard coordinate functions on the $2$-dimensional summand, and $t_{1},t_{2}$ stand for the coordinate functions on the $1$-dimensional summands. The invariant $(x_{1}^{n}-x_{2}^{n})t_{1}t_{2}$ separates the points $v:=([1,0]^{T},1,1)$ and $v^{\prime}:=([1,0]^{T},1,-1)$ in $K^{4}$. It is sufficient to show that $v$ and $v^{\prime}$ can not be separated by an invariant of degree at most $n+1$. Suppose for contradiction that $h(v)\neq h(v^{\prime})$ for some homogeneous $h\in K[x_{1},x_{2},t_{1},t_{2}]^{G}$ of degree at most $n+1$. We may assume by Lemma 3.5 that $h$ is multihomogeneous (so it is homogeneous both in $t_{1}$ and $t_{2}$), and has minimal possible total degree. So $h=t_{1}^{k_{1}}t_{2}^{k_{2}}h_{1}(x_{1},x_{2})$. Since the $(x_{1},x_{2},t_{1})$-coordinates of $v$ and $v^{\prime}$ are the same, we have $k_{2}>0$. Since $h_{1}$ is $\langle c\rangle$-invariant, $k_{1}+k_{2}$ must be even. Moreover, $K[t_{1},t_{2}]^{G}=K[t_{1}^{2},t_{2}^{2}]$, and $t_{1}^{2}(v)=1=t_{1}^{2}(v^{\prime})$, $t_{2}^{2}(v)=1=t_{2}^{2}(v^{\prime})$, so $\deg(h_{1})>0$, and (since $h$ has minimal possible degree) $h=t_{1}t_{2}h_{1}$. Thus $h_{1}$ is a relative $G$-invariant of degree at most $n-1$, on which $G$ acts via the character $a\mapsto 1,$ $b\mapsto-1$, $c\mapsto 1$. The action of $G$ on $K^{2}$ factors through the action of the dihedral group $\mathrm{D}_{2n}\cong G/\langle c\rangle$, and it is well known (and easy to see) that the smallest degree relative invariant in $K[x_{1},x_{2}]$ with the above weight is $x_{1}^{n}-x_{2}^{n}$. We reached the desired contradiction. ∎
Proof of Theorem 2.2. We have the isomorphism
$\mathrm{D}_{2n}\times\mathrm{C}_{2}\cong(\mathrm{C}_{n}\times\mathrm{C}_{2})\rtimes_{-1}\mathrm{C}_{2}.$
Therefore by [12, Corollary 5.5] we have
$\beta^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})=\mathsf{D}(\mathrm{C}_{n}\times\mathrm{C}_{2})+1=n+2,$
implying in turn that $\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})\leq n+2$. The reverse inequality holds by Proposition 4.3. $\square$
5. Further groups $G$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta^{K}(G)$
5.1. The alternating group $\mathrm{A}_{4}$
Proposition 5.1.
For $n\geq 4$ we have the inequality $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{n})\geq n(n-1)/2$.
Proof.
Our running assumption that $\mathrm{char}(K)$ does not divide $|G|$ forces for $G=\mathrm{A}_{n}$ that $|K|\geq n$ if $n\geq 4$. Therefore $K$ has $n$ distinct elements $v_{1},\dots,v_{n}$. The vectors $v:=[v_{1},v_{2},v_{3},\dots,v_{n}]^{T}$ and $v^{\prime}:=[v_{2},v_{1},v_{3},v_{4},\dots,v_{n}]^{T}$ belong to the same orbit under the symmetric group $\mathrm{S}_{n}$ acting on $K^{n}$ via permuting the coordinates, but $v$ and $v^{\prime}$ have different orbits with respect to the alternating subgroup $\mathrm{A}_{n}$. It is well known that the corresponding algebra of invariants is
$K[x_{1},\dots,x_{n}]^{\mathrm{A}_{n}}=K[x_{1},\dots,x_{n}]^{\mathrm{S}_{n}}\oplus\Delta K[x_{1},\dots,x_{n}]^{\mathrm{S}_{n}},$
where $\Delta:=\sum_{1\leq i<j\leq n}(x_{i}-x_{j})$. This shows that all $\mathrm{A}_{n}$-invariants of degree less than $n(n-1)/2$ are in fact $\mathrm{S}_{n}$-invariants. So they can not separate $v$ and $v^{\prime}$, which have different $\mathrm{A}_{n}$-orbits. ∎
It is known that $\beta^{K}(\mathrm{A}_{4})=6$ (see [11, Theorem 3.4]), hence Proposition 5.1 has the following consequence:
Theorem 5.2.
We have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})=6$.
@@ 5.2. The binary tetrahedral group
In this section $G:=\widetilde{\mathrm{A}}_{4}$ is the binary tetrahedral group: the special orthogonal group $\mathrm{SO}(3,\mathbb{R})$ has a subgroup isomorphic to $\mathrm{A}_{4}$ (the group of rotations mapping a regular tetrahedron into itself), and $G$ is isomorphic to the preimage of $\mathrm{A}_{4}$ under the classical two-fold covering $\mathrm{SU}(2,\mathbb{C})\to\mathrm{SO}(3,\mathbb{R})$. So $G$ can be embedded as a subgroup into the special unitary group $\mathrm{SU}(2,\mathbb{C})$, acting via its defining representation on $V:=\mathbb{C}^{2}$. Thus $V$ is a $G$-module, and it is well known that $\mathbb{C}[V]^{G}$ is generated by three homogeneous invariants $f_{6}$, $f_{8}$, $f_{12}$, having degree $6$, $8$, and $12$ (see [32, Lemma 4.1] or [42, Section 0.13]). Since $G$ (as a subgroup of $\mathrm{GL}(V)$) is not generated by pseudo-reflections, $\mathbb{C}[V]^{G}$ does not contain an algebraically independent separating set by [24], so the invariants $f_{6}$ and $f_{8}$ can not separate all the $G$-orbits in $V$. This implies that $\beta_{\mathrm{sep}}^{\mathbb{C}}(G,V)\geq 12$. A second proof valid for more general base fields is given below.

### v3

@@ 3.1. The group $(\mathrm{C}_{3}\times\mathrm{C}_{3})\rtimes_{-1}\mathrm{C}_{2}$
Note that $\ker(\psi_{1})=\langle b\rangle$ and $\ker(\psi_{2})=\langle a\rangle$. We have $G/\ker(\psi_{1})\cong\mathrm{S}_{3}$ and $G/\ker(\psi_{2})\cong\mathrm{S}_{3}$. Both these representations factor through the irreducible $2$-dimensional representation of $\mathrm{S}_{3}$. It is well known and easy to see that the corresponding algebra of invariants is generated by a degree $2$ and a degree $3$ invariant. Moreover, $K[W_{j}]/\mathcal{H}(G,W_{j})$ as an $\mathrm{S}_{3}=G/\ker(\psi_{j})$-module is isomorphic to the regular representation of $\mathrm{S}_{3}$. Thus $x_{1}^{3}-x_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $KG$-module direct complement of $\mathcal{H}(G,W_{1})$ in $K[W_{1}]$ on which $\ker(\psi_{2})$ acts trivially, and $y_{1}^{3}-y_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $KG$-module direct complement of $\mathcal{H}(G,W_{2})$ in $K[W_{2}]$ on which $\ker(\psi_{1})$ acts trivially. It follows that $K[W_{1}\oplus W_{2}]^{G}$ is generated by $K[W_{1}]^{G}$, $K[W_{2}]^{G}$, and $f:=(x_{1}^{3}-x_{2}^{3})(y_{1}^{3}-y_{2}^{3})$. Consider the points $v=(w_{1},w_{2})=([1,0]^{T},[1,0]^{T})$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2})=([1,0]^{T},[0,1]^{T})$. We claim that all invariants of degree at most $5$ agree on $v$ and on $v^{\prime}$, but $f(v)\neq f(v^{\prime})$. Indeed, $w_{1}=w^{\prime}_{1}$ and $c\cdot w_{2}=w^{\prime}_{2}$. Thus the invariants in $K[W_{1}]^{G}$ and in $K[W_{2}]^{G}$ agree on $v$ and on $v^{\prime}$. On the other hand, $f(v)=1$ and $f(v^{\prime})=-1$. The proof is finished. ∎
Theorem 3.2.
Proof.
By [8, Corollary 5.5] we have $\beta^{K}(G)=6$. Thus the result follows from Proposition 3.1. ∎
3.2. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
The value of $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{7}\rtimes\mathrm{C}_{3})$ is obtained as a special case of a result in [5] under the assumption that $K$ is algebraically closed. This can be generalized to the case when $K$ contains an element of multiplicative order $|G|$.
@@ 3.2. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
[5, Theorem 4.1] Suppose that $p\equiv 1\mod 3$ is a prime and $K$ is algebraically closed. Then for the group $\mathrm{C}_{p}\rtimes\mathrm{C}_{3}=\langle a,b\mid a^{p}=b^{3}=1_{G},\quad bab^{-1}=a^{2}\rangle$ we have $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{p}\rtimes\mathrm{C}_{3})=p+1$.
Corollary 3.4.
Consider the group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}=\langle a,b\mid a^{7}=b^{3}=1_{G},\quad bab^{-1}=a^{2}\rangle$ and suppose that $K$ contains an element of multiplicative order $21$. Then $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{7}\rtimes\mathrm{C}_{3})=8$.
@@
Actually, [5] works over an algebraically closed base field, but the representation providing the exact lower bound for the separating Noether number is defined over any field containing an element of multiplicative order $3p$, and in order to prove that $p+1$ is also an upper bound, we may pass to the algebraic closure using Lemma 2.3. As the special case $p=7$, we are done. ∎
@@ 3.3. The binary tetrahedral group
In this subsection $G:=\widetilde{\mathrm{A}}_{4}$ is the binary tetrahedral group: the special orthogonal group $\mathrm{SO}(3,\mathbb{R})$ has a subgroup isomorphic to $\mathrm{A}_{4}$ (the group of rotations mapping a regular tetrahedron into itself), and $G$ is isomorphic to the preimage of $\mathrm{A}_{4}$ under the classical two-fold covering $\mathrm{SU}(2,\mathbb{C})\to\mathrm{SO}(3,\mathbb{R})$. So $G$ can be embedded as a subgroup into the special unitary group $\mathrm{SU}(2,\mathbb{C})$, acting via its defining representation on $V:=\mathbb{C}^{2}$. Thus $V$ is a $KG$-module, and it is well known that $\mathbb{C}[V]^{G}$ is generated by three homogeneous invariants $f_{6}$, $f_{8}$, $f_{12}$, having degree $6$, 88, and $12$ (see [21, Lemma 4.1] or [29, Section 0.13]). Since $G$ (as a subgroup of $\mathrm{GL}(V)$) is not generated by pseudo-reflections, $\mathbb{C}[V]^{G}$ does not contain an algebraically independent separating set by [17], so the invariants $f_{6}$ and $f_{8}$ can not separate all the $G$-orbits in $V$. This implies that $\beta_{\mathrm{sep}}^{\mathbb{C}}(G,V)\geq 12$. A second proof valid for more general base fields is given below.